OpenStudy (anonymous):

How do I write the expression (-27)^5/3 in radical notation How do I evaluate the expression?

OpenStudy (anonymous):

is it \[(-27)^{5/3}\]

OpenStudy (anonymous):

using the law (a)^m.n = (a^m)^n \[(-27)^{5/3} = [(-3)^{3}]^{5/3} = (-3)^{3*5/3} = (-3)^{5} = -243\]

OpenStudy (anonymous):

thank u.....